### v1

Abstract
Adding additional control to pretrained diffusion models has become an increasingly popular research area, with extensive applications in computer vision, reinforcement learning, and AI for science. Recently, several studies have proposed training-free diffusion guidance by using off-the-shelf networks pretrained on clean images. This approach enables zero-shot conditional generation for universal control formats, which appears to offer a free lunch in diffusion guidance. In this paper, we aim to develop a deeper understanding of the operational mechanisms and fundamental limitations of training-free guidance. We offer a theoretical analysis that supports training-free guidance from the perspective of optimization, distinguishing it from classifier-based (or classifier-free) guidance. To elucidate their drawbacks, we theoretically demonstrate that training-free methods are more susceptible to adversarial gradients and exhibit slower convergence rates compared to classifier guidance. We then introduce a collection of techniques designed to overcome the limitations, accompanied by theoretical rationale and empirical evidence. Our experiments in image and motion generation confirm the efficacy of these techniques.
1 Introduction
1.1 Motivations
Diffusion models represent a class of powerful deep generative models that have recently broken the long-standing dominance of generative adversarial networks (GANs) [7]. These models have demonstrated remarkable success in a variety of domains, including the generation of images and videos in computer vision [29, 3], the synthesis of molecules and proteins in computational biology [16, 47], as well as the creation of trajectories and actions in the field of reinforcement learning (RL) [17].
One critical area of research in the field of diffusion models involves enhancing controllability, such as pose manipulation in image diffusion [45], modulation of quantum properties in molecule diffusion [16], and direction of goal-oriented actions in RL diffusion [19]. The predominant techniques for exerting control over diffusion models include classifier guidance and classifier-free guidance. Classifier guidance involves training a time-dependent classifier to map a noisy image, denoted as $\bm{x}_{t}$, to a specific condition $\bm{y}$, and then employing the classifier’s gradient to influence each step of the diffusion process [7]. Conversely, classifier-free guidance bypasses the need for a classifier by training an additional diffusion model conditioned on $\bm{y}$ [15]. However, both approaches necessitate extra training to integrate the conditions. Moreover, their efficacy is often constrained when the data-condition pairs are limited and typically lacks the capability for zero-shot generalization to novel conditions.
Recently, several studies [2, 44, 33] have introduced training-free guidance that builds upon the concept of classifier guidance. These models eschew the need for training a classifier on noisy images; instead, they estimate the clean image from its noisy counterpart using Tweedie’s formula and then employ pre-trained networks, designed for clean images, to guide the diffusion process. Given that checkpoints for these networks pretrained on clean images are widely accessible online, this form of guidance can be executed in a zero-shot manner. A unique advantage of training-free guidance is that it can be applied to universal control formats, such as style, layout, and FaceID [2, 44, 33] without any additional training efforts. Furthermore, these algorithms have been successfully applied to offline reinforcement learning, enabling agents to achieve novel goals not previously encountered during training [40]. Despite the encouraging empirical results, the fundamental principles contributing to the success of these methods have yet to be fully elucidated. From a theoretical perspective, it is intriguing to understand how and when these methods succeed or fail. From an empirical standpoint, it is crucial to develop algorithms that can address and overcome these limitations.
This paper seeks to deepen the understanding of training-free guidance by examining its operational mechanisms and inherent limitations. From an optimization standpoint, we demonstrate that training-free guidance has a two-stage convergence, where the second stage enjoys a linear convergence. This is a distinctive feature that is not shared by classifier guidance or classifier-free guidance methods. However, we identify and provide proof that training-free guidance is prone to adversarial gradient issues and slower convergence rates due to the diminished smoothness of the guidance network, when compared to classifier guidance. To address these challenges, we introduce a suite of techniques that have been theoretically and empirically validated for their efficacy. Specifically, our major contributions can be summarized as follows:
  • Mechanisms (How it works): We offer theoretical proof, from an optimization perspective, that training-free guidance is guaranteed to minimize the loss of the guidance network. This characteristic distinguishes them from approaches that rely on training-based guidance.
  • Limitations (When it does not work): We theoretically identify the susceptibility of training-free guidance to adversarial gradient issues and slower convergence rates. We attribute these challenges to a decrease in the smoothness of the guidance network in contrast to the classifier guidance.
  • Addressing the limitations: We introduce a set of enhancement techniques aimed at mitigating the identified limitations. The efficacy of these methods is empirically confirmed across various diffusion models (e.g., image diffusion and motion diffusion) and under multiple conditions (e.g., segmentation, sketch, text, and object avoidance). The code is available at https://github.com/BIGKnight/Understanding-Training-free-Diffusion-Guidance.
For more related fields and the impact of this paper in these fields, please refer to Appendix 7.
2 Preliminaries
2.1 Diffusion Models
Diffusion models are characterized by forward and reverse processes. The forward process, occurring over a time interval from 00 to $T$, incrementally transforms an image into Gaussian noise. On the contrary, the reverse process, from $T$ back to 00, reconstructs the image from the noise. Let $\bm{x}_{t}$ represent the state of the data point at time $t$; the forward process systematically introduces noise to the data by following a predefined noise schedule given by
$\bm{x}_{t}=\sqrt{\alpha_{t}}\bm{x}_{0}+\sigma_{t}\bm{\epsilon}_{t},$
where $\alpha_{t}\in[0,1]$ is monotonically decreasing with $t$, $\sigma_{t}=\sqrt{1-\alpha_{t}}$, and $\bm{\epsilon}_{t}\sim\mathcal{N}(0,\bm{I})$ is random noise. Diffusion models use a neural network to learn the noise at each step:
$\min_{\theta}\mathbb{E}_{\bm{x}_{t},\bm{\epsilon},t}[\|\bm{\epsilon}_{\theta}(\bm{x}_{t},t)-\bm{\epsilon}_{t}\|_{2}^{2}]=\min_{\theta}\mathbb{E}_{\bm{x}_{t},\bm{\epsilon},t}[\|\bm{\epsilon}_{\theta}(\bm{x}_{t},t)+\sigma_{t}\nabla_{\bm{x}_{t}}\log p_{t}(\bm{x}_{t})\|_{2}^{2}],$
where $p_{t}(\bm{x}_{t})$ is the distribution of $\bm{x}_{t}$. The reverse process is obtained by the following ODE:
$\frac{{\rm d}\bm{x}_{t}}{{\rm d}t}$ $=f(t)\bm{x}_{t}-\frac{g^{2}(t)}{2}\nabla_{\bm{x}_{t}}\log p_{t}(\bm{x}_{t})=f(t)\bm{x}_{t}+\frac{g^{2}(t)}{2\sigma_{t}}\bm{\epsilon}_{\theta}(\bm{x}_{t},t),$ (1)
where $f(t)=\frac{{\rm d}\log\sqrt{\alpha_{t}}}{{\rm d}t}$, $g^{2}(t)=\frac{{\rm d}\sigma_{t}^{2}}{{\rm d}t}-2\frac{{\rm d}\log\sqrt{\alpha_{t}}}{{\rm d}t}\sigma_{t}^{2}$. The reverse process enables generation as it converts a Gaussian noise into the image.
2.2 Training-Free Diffusion Guidance
In diffusion control, we aim to sample $\bm{x}_{0}$ given a condition $\bm{y}$. The conditional score function is expressed as follows:
$\nabla_{\bm{x}_{t}}\log p_{t}(\bm{x}_{t}|\bm{y})=\nabla_{\bm{x}_{t}}\log p_{t}(\bm{x}_{t})+\nabla_{\bm{x}_{t}}\log p_{t}(\bm{y}|\bm{x}_{t}).$ (2)
The conditions are specified by the output of a neural network, and the energy is quantified by the corresponding loss function. If $\ell(f_{\phi}(\cdot),\cdot)$ represents the loss function as computed by neural networks, then the distribution of the clean data is expected to follow the following formula [2, 44, 33]:
$p_{0}(\bm{x}_{0}|\bm{y})\propto p_{0}(\bm{x}_{0})\exp(-\ell(f_{\phi}(\bm{x}_{0}),\bm{y})).$ (3)
For instance, consider a scenario where the condition is the object location. In this case, $f_{\phi}$ represents a fastRCNN architecture, and $\ell$ denotes the classification loss and bounding box loss. By following the computations outlined in [22], we can derive the exact formula for the second term in the RHS of (2) as:
$\nabla_{\bm{x}_{t}}\log p_{t}(\bm{y}|\bm{x}_{t})=\nabla_{\bm{x}_{t}}\log\mathbb{E}_{p(\bm{x}_{0}|\bm{x}_{t})}[\exp(-\ell(f_{\phi}(\bm{x}_{0}),\bm{y})].$ (4)
Classifier guidance [7] involves initially training a time-dependent classifier to predict the output of the clean image $\bm{x}_{0}$ based on noisy intermediate representations $\bm{x}_{t}$ during the diffusion process, i.e., to train a time-dependent classifier $f_{\psi}(\bm{x}_{t},t)$ such that $f_{\psi}(\bm{x}_{t},t)\approx f_{\phi}(\bm{x}_{0})$. Then the gradient of the time-dependent classifier is used for guidance, given by $\nabla_{\bm{x}_{t}}\log p_{t}(\bm{y}|\bm{x}_{t}):=-\nabla_{\bm{x}_{t}}\ell(f_{\psi}(\bm{x}_{t},t),\bm{y})$.
Classifier guidance corresponds to the energy function in (4) when the loss function employed is cross-entropy. This direct relationship does not hold when alternative loss functions, such as the $\ell_{2}$ distance [22], are utilized. Nonetheless, empirical research [20] indicates that classifier guidance remains effective across a diverse array of applications, including but not limited to text guidance and image similarity guidance.
Training-free guidance [2, 44, 6] puts the expectation in (4) inside the loss function:
$\nabla_{\bm{x}_{t}}\log p_{t}(\bm{y}|\bm{x}_{t}):=$ $\nabla_{\bm{x}_{t}}\log\left[\exp(-\ell(f_{\phi}(\mathbb{E}_{p(\bm{x}_{0}|\bm{x}_{t})}(\bm{x}_{0})),\bm{y})\right]$ (5)
=$\overset{(a)}{=}$ $-\nabla_{\bm{x}_{t}}\ell\left[f_{\phi}\left(\frac{\bm{x}_{t}-\sigma_{t}\bm{\epsilon}_{\theta}(\bm{x}_{t},t)}{\sqrt{\alpha_{t}}}\right),\bm{y}\right],$
where (a) uses Tweedie’s formula $\mathbb{E}_{p(\bm{x}_{0}|\bm{x}_{t})}(\bm{x}_{0})=\frac{\bm{x}_{t}-\sigma_{t}\bm{\epsilon}_{\theta}(\bm{x}_{t},t)}{\sqrt{\alpha_{t}}}$. Leveraging this approximation permits the use of a pre-trained off-the-shelf network designed for processing clean data. The gradient of the last term in the energy function is obtained via backpropagation through both the guidance network and the diffusion backbone.
3 Analysis of Training-Free Guidance
3.1 Mechanisms of Training-free Guidance
3.1.1 The difficulty of approximating $p_{t}(\bm{y}|\bm{x}_{t})$ in high-dimensions.
From a probabilistic standpoint, it is observed that [33]
$\nabla_{\bm{x}_{t}}\log\mathbb{E}_{p(\bm{x}_{0}|\bm{x}_{t})}[\exp(-\ell(f_{\phi}(\bm{x}_{0}),\bm{y})]\overset{(a)}{\approx}\nabla_{\bm{x}_{t}}\log\mathbb{E}_{q(\bm{x}_{0}|\bm{x}_{t})}[\exp(-\ell(f_{\phi}(\bm{x}_{0}),\bm{y})]$ (6)
$\approx\nabla_{\bm{x}_{t}}\log\frac{1}{n}\sum_{i=1}^{n}\exp(-\ell(f_{\phi}(\bm{x}^{i}_{0}),\bm{y}),\quad\bm{x}^{i}_{0}\sim q(\bm{x}_{0}|\bm{x}_{t}),$
where $q(\bm{x}_{0}|\bm{x}_{t})=\mathcal{N}(\mathbb{E}_{p(\bm{x}_{0}|\bm{x}_{t})}(\bm{x}_{0}),r_{t}^{2}\bm{I})$ and $r_{t}$ is a tunable parameter [33]. By taking $n=1$ and $r_{t}=0$, (6) is reduced to the guidance term in (5). As demonstrated in [33], the approximation is effective for one-dimensional distribution. However, the approximation denoted by (a) does not extend to high-dimensional data, such as images. This is due to the well-known high-dimensional probability phenomenon where $q(\bm{x}_{0}|\bm{x}_{t})$ tends to concentrate on a spherical shell centered at $\mathbb{E}_{p(\bm{x}_{0}|\bm{x}_{t})}(\bm{x}_{0})$ with radius $r_{t}$, as discussed in Chapter 3.1 of [39]. Since the spherical shell represents a low-dimensional manifold with zero measure in the high-dimensional space, there is a significant likelihood that the supports of $p(\bm{x}_{0}|\bm{x}_{t})$ and $q(\bm{x}_{0}|\bm{x}_{t})$ do not overlap, rendering the approximation (a) ineffective.
3.1.2 Training-free guidance minimizes the loss from an optimization perspective.
Given the difficulty of analyzing the training-free guidance from a probability perspective, we turn to an optimization perspective for clarity. Denote $\hat{\bm{x}}_{0}=\mathbb{E}_{p(\bm{x}_{0}|\bm{x}_{t})}(\bm{x}_{0})$, the gradient guidance term in (5) can be written as the following:
$\nabla_{\bm{x}_{t}}\ell\left[f_{\phi}(\hat{\bm{x}}_{0}),\bm{y}\right]=\frac{\partial\ell}{\partial\hat{\bm{x}}_{0}}\nabla_{\bm{x}_{t}}\left(\frac{\bm{x}_{t}+\sigma_{t}^{2}\nabla_{x_{t}}\log p_{t}(\bm{x}_{t})}{\sqrt{\alpha_{t}}}\right)=\frac{\partial\ell}{\partial\hat{\bm{x}}_{0}}\frac{\text{Cov}[\bm{x}_{0}|\bm{x}_{t}]}{\sigma_{t}^{2}\sqrt{\alpha_{t}}},$ (7)
where the last equality follows the second-order Miyasawa relationship $\text{Cov}[\bm{x}_{0}|\bm{x}_{t}]=\sigma_{t}^{2}(\bm{I}+\sigma_{t}^{2}\nabla^{2}\log p_{t}(\bm{x}_{t}))$ [26]. Based on (7), we show that the guidance steps decrease the loss function associated with the clean classifier $f_{\phi}$, which is shown in the next proposition.
Proposition 1
Assume that the guidance loss function $\ell(f_{\phi}(\bm{x}_{0}),\bm{y})$ is $\mu$-PL (defined in Defintion 2 in appendix) and $L_{f}$-Lipschitz with respect to clean images $\bm{x}_{0}$, and the score function $\nabla\log p_{t}(\bm{x}_{t})$ is $L_{p}$-Lipschitz (defined in Defintion 1 in appendix) with respect to noisy image $\bm{x}_{t}$. Then the following conditions hold:
  1. 1.
[…]
  2. 2.
    Denote $\kappa_{1}=\frac{\mu\lambda_{\min}^{2}}{L_{f}(1+L_{p})\sqrt{\alpha_{t}}\sigma_{t}^{4}}$, After one gradient step $\hat{\bm{x}}_{t}=\bm{x}_{t}-\eta\nabla_{\bm{x}_{t}}\ell_{t}(\bm{x}_{t}),\eta=\frac{\sqrt{\alpha_{t}}}{L_{f}(1+L_{p})}$, we have $\ell_{t}(\hat{\bm{x}}_{t})\leq(1-\kappa_{1})\ell_{t}(\bm{x}_{t})$;
  3. 3.
    Consider a diffusion process that adheres to a bounded change in the objective function such that for any diffusion step, i.e., $\ell_{t-1}(\bm{x}_{t-1})\leq\frac{\ell_{t}(\hat{\bm{x}}_{t})}{(1-\kappa_{2})}$ for some $\kappa_{2}<\kappa_{1}$, then the objective function converges at a linear rate, i.e., $\ell_{t-1}(\bm{x}_{t-1})\leq\frac{1-\kappa_{1}}{1-\kappa_{2}}\ell_{t}(\bm{x}_{t})$.
(a) A successful guidance.
(b) A failure guidance.
Figure 1: The classifier loss of a successful and a failure guidance example. The target class is “indigo bird”.
The proof is given in Appendix 8.2. The Lipschitz continuity and PL conditions are basic assumptions in optimization, and it has been shown that neural networks can locally satisfy these conditions [5]. Our next focus is on scenarios where the assumption in the third condition is met. Specifically, for large values of diffusion time $t$, the diffusion trajectory can exhibit substantial deviations between adjacent steps, leading to an increased value of $\kappa_{2}$ in the third condition. This phenomenon is less pronounced for smaller $t$. Consequently, the theorem leads us to anticipate a two-phase convergence pattern in training-free guidance: an initial phase where the loss remains relatively high and oscillates for larger $t$, followed by a swift decline (linear convergence) towards lower loss values as $t$ decreases. In Figure 1, we use ResNet-50 trained on clean images to guide ImageNet pre-trained diffusion models. The loss value at each diffusion step is plotted. As a reference, we choose $100$ images from the class “indigo bird” in ImageNet training set and compute the loss value, which is referred as “Correct Image Loss” in the figure. The two-stage convergence is verified by experiments. Remarkably, from theory and experiments, we see that training-free guidance can achieve low loss levels even in instances of guidance failure (Figure 0(b)). This is attributed to the influence of adversarial gradients, as discussed in the next subsection.
3.2 Limitations of Training-free Guidance
In this subsection, we examine the disadvantages of employing training-free guidance networks as opposed to training-based classifier guidance.
3.2.1 Training-free guidance is more sensitive to the adversarial gradient.
Adversarial gradient is a significant challenge for neural networks, which refers to minimal perturbations deliberately applied to inputs that can induce disproportionate alterations in the model’s output [34]. The resilience of a model to adversarial gradients is often analyzed through the lens of its Lipschitz constant [31]. If the model has a lower Lipschitz constant, then it is less sensitive to the input perturbations and thus is more robust. In classifier guidance, time-dependent classifiers are trained on noise-augmented images. Our finding is that adding Gaussian noise transforms non-Lipschitz classifiers into Lipschitz ones. This transition mitigates the adversarial gradient challenge by inherently enhancing the model’s robustness to such perturbations, as shown in the next proposition.
Proposition 2
(Time-dependent network is more robust and smooth) Given a bounded non-Lipschitz loss function $\ell(\bm{x})\leq C$, the loss ^$\hat{\ell}(\bm{x})=\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(0,\bm{I})}[\ell(\bm{x}+\sigma_{t}\bm{\epsilon})]$ is $C\sqrt{\frac{2}{\pi\sigma_{t}^{2}}}$-Lipschitz and $\nabla\hat{\ell}$ is $\frac{2C}{\sigma_{t}}$-Lipschitz.
(a) Adversarially robust classifier.
(b) Time-dependent classifier.
(c) Off-the-shelf ResNet-50 classifier.
(d) ResNet-50 with random augmentation.
Figure 2: Gradients of different classifiers on random backgrounds. The images in the first row correspond to the target class “cock”, and the second row to “goldfinch”.
The proof is given in Appendix 8.3. We present visualizations of the accumulated gradients for both the time-dependent and off-the-shelf time-independent classifiers corresponding to different classes in Figure 1(b) and Figure 1(c), respectively. These visualizations are generated by initializing an image with a random background and computing $1000$ gradient steps for each classifier. For the time-dependent classifier, the input time for the $t$-th gradient step is $1000-t$. The images are generated purely by the classifier gradients without diffusion involved. For comparative analysis, we include the accumulated gradient of an adversarially robust classifier [30], as shown in Figure 1(a), which has been specifically trained to resist adversarial gradients. The resulting plots reveal a stark contrast: the gradient of the time-dependent classifier visually resembles the target image, whereas the gradient of the time-independent classifier does not exhibit such recognizability. This observation suggests that off-the-shelf time-independent classifiers are prone to generating adversarial gradients compared to the time-dependent classifier used in classifier guidance. It could potentially misdirect the guidance process. In contrast to yielding a direction that meaningfully minimizes the loss, the adversarial gradient primarily serves to minimize the loss in a manner that is not necessarily aligned with the intended guidance direction.
3.2.2 Training-free guidance slows down the convergence of reverse ODE.
The efficiency of an algorithm in solving reverse ordinary differential equations is often gauged by the number of non-linear function estimations (NFEs) required to achieve convergence. This metric is vital for algorithmic design, as it directly relates to computational cost and time efficiency [32]. In light of this, we explore the convergence rates associated with various guidance paradigms, beginning our analysis with a reverse ODE framework that incorporates a generic gradient guidance term. The formula is expressed as
$\frac{{\rm d}\bm{x}_{t}}{{\rm d}t}$ $=f(t)\bm{x}_{t}+\frac{g^{2}(t)}{2}(\bm{\epsilon}_{\theta}(\bm{x}_{t},t)+\nabla_{\bm{x}_{t}}h(\bm{x}_{t},t)),$ (8)
where $h(\cdot,\cdot)$ can be either a time-dependent classifier or a time-independent classifier with Tweedie’s formula. The subsequent proposition elucidates the relationship between the discretization error and the smoothness of the guidance function.
Proposition 3
Let $g(\bm{x}_{t},t)=\bm{\epsilon}_{\theta}(\bm{x}_{t},t)+\nabla_{\bm{x}_{t}}h(\bm{x}_{t},t)$ in (8). Assume that $g(\bm{x}_{t},t)$ is $L$-Lipschitz. Then the discretization error of the DDIM solver is bounded by $O(h_{\max}+Lh_{\max})$, where $h_{\max}=\max_{t}\frac{1}{2}\left[\log(\frac{\alpha_{t}}{1-\alpha_{t}})-\log(\frac{\alpha_{t-1}}{1-\alpha_{t-1}})\right]$.
The proof is given in Appendix 8.4. Proposition 2 establishes that time-dependent classifiers exhibit superior gradient Lipschitz constants compared to their off-the-shelf time-independent counterparts. This disparity in smoothness attributes leads to a deceleration in convergence for training-free guidance methods, necessitating a greater number of NFEs to achieve the desired level of accuracy when compared to classifier guidance.
4 Improving Training-free Guidance
This section presents enhancement techniques for improving training-free guidance, with theoretical justifications and empirical evidence.
Algorithm 1 Random Augmentation
for $t=T,\cdots,0$ do
     $\bm{x}_{t-1}=\text{DDIM}(\bm{x}_{t})$
     $\hat{\bm{x}}_{0}=\frac{\bm{x}_{t}-\sigma_{t}\bm{\epsilon}_{\theta}(\bm{x}_{t},t)}{\sqrt{\alpha_{t}}}$
     $\bm{g}_{t}=\frac{1}{|\mathcal{T}|}\sum_{T\in\mathcal{T}}\nabla_{\bm{x}_{t}}\ell(f_{\phi}(T(\hat{\bm{x}}_{0})),\bm{y})$
     $\bm{x}_{t-1}=\bm{x}_{t-1}-\eta\bm{g}_{t}$
end for
Algorithm 2 PGD
for $t=T,\cdots,0$ do
     $\bm{x}_{t-1}=\text{DDIM}(\bm{x}_{t})$
     $\hat{\bm{x}}_{0}=\frac{\bm{x}_{t}-\sigma_{t}\bm{\epsilon}_{\theta}(\bm{x}_{t},t)}{\sqrt{\alpha_{t}}}$
     $\bm{g}_{t}=\nabla_{\bm{x}_{t}}\ell(f_{\phi}(\hat{\bm{x}}_{0}),\bm{y})$
     $\bm{x}_{t-1}=\bm{x}_{t-1}-\eta\cdot\frac{\bm{g}_{t}}{\|\bm{g}_{t}\|_{2}}$
end for
4.1 Randomized Augmentation
As established by Proposition 2, the introduction of Gaussian perturbations enhances the Lipschitz property of a neural network. A direct application of this principle involves creating multiple noisy instances of an estimated clean image and incorporating them into the guidance network, a method analogous to the one described in (6). However, given the high-dimensional nature of image data, achieving a satisfactory approximation of the expected value necessitates an impractically large number of noisy copies. To circumvent this issue, we propose an alternative strategy that employs a diverse set of data augmentations in place of solely adding Gaussian noise. This approach effectively introduces perturbations within a lower-dimensional latent space, thus requiring fewer samples. The suite of data augmentations utilized, denoted by $\mathcal{T}$, is derived from the differentiable data augmentation techniques outlined in [46], which encompasses transformations such as translation, resizing, color adjustments, and cutout operations. The details are shown in Algorithm 1 and the rationale is shown in the following proposition.
Proposition 4
(Random augmentation improves smoothness) Given a bounded non-Lipschitz loss function $\ell(\bm{x})$, the loss ^$\hat{\ell}(\bm{x})=\mathbb{E}_{\bm{\epsilon}\sim p(\bm{\epsilon})}[\ell(\bm{x}+\bm{\epsilon})]$ is
[…]
The proof is shown in Appendix 8.5. Echoing the experimental methodology delineated in Section 3.2, we present an analysis of the accumulated gradient effects when applying random augmentation to a ResNet-50 model. Specifically, we utilize a set of $|\mathcal{T}|=10$ diverse transformations as our augmentation strategy. The results of this experiment are visualized in Figure 1(d), where the target object’s color and shape emerges in the gradient profile. This observation suggests that the implementation of random augmentation can alleviate the adversarial gradient issue.
4.2 Adaptive Gradient Scheduling
In Section 3.1, we established that the training guidance process can be conceptualized as the optimization of the loss function corresponding to the guidance network. Given this interpretation, it stands to reason that adopting a more sophisticated optimizer could accelerate the convergence of the guidance. Notably, the fundamental operation in guidance involves calculating the gradient of the guidance network with respect to the input, a procedure reminiscent of white-box adversarial attacks. Drawing inspiration from this parallel, we adopt the well-known Projected Gradient Descent (PGD) optimization technique—commonly employed in crafting adversarial examples—as our optimizer of choice, which is shown in Algorithm 2. We implement PGD within the context of a training-free guidance framework called FreeDoM [44] and benchmark the performance of this implementation using the DDIM sampler with $50$ steps, ensuring identical seed initialization across the different methods for consistency.
As shown in Figure 3, FreeDoM is unable to effectively guide the generation process when faced with a significant discrepancy between the unconditional generation and the specified condition. An illustrative example is the difficulty in guiding the model to generate faces oriented to the left when the unconditionally generated faces predominantly orient to the right, as shown in the middle column of the figure. Conversely, when the unconditionally generated faces orient forward, the condition for any face orientation is more easily fulfilled. This challenge, which arises due to the insufficiency of $50$ steps for convergence under the condition, is ameliorated by substituting gradient descent with PGD, thereby illustrating the benefits of employing an advanced optimizer in the guidance process.
Figure 3: FreeDoM [44] with and without PGD trick. DDIM-$50$ is used for sampling.
4.3 Resampling Trick
Algorithm 3 Resampling
1:for $t=T,\cdots,1$ do
2:     for $i=1,\cdots,s$ do
3:         $\bm{x}_{t-1}^{i}=\text{DDIM with Guidance}(\bm{x}_{t}^{i-1})$
4:         if $i<s$ then
5:              $\beta_{t}=\alpha_{t}/\alpha_{t-1},\bm{n}\sim\mathcal{N}(0,\bm{I})$
6:              $\bm{x}_{t}^{i}=\sqrt{\beta_{t}}\bm{x}_{t-1}^{i}+\sqrt{1-\beta_{t}}\bm{n}$
7:         end if
8:         $\bm{x}_{t-1}^{0}=\bm{x}_{t-1}^{s}$
9:     end for
10:end for
The technique of resampling, also referred to as “time travel”, has been proposed as a solution to complex generative problems [24], facilitating successful training-free guidance in tasks such as CLIP-guided ImageNet generation and layout guidance for stable diffusion, as illustrated in Figure 2 of [44] and Figure 8 of [2], respectively. The procedure of the resampling trick is shown in Algorithm 3, which involves recursive execution of individual sampling steps.
In this subsection, we endeavor to analyze the role of resampling in training-free guidance and thus complete the overall picture. During the diffusion process, samples can diverge from their target distributions due to various types of sampling error, such as adversarial gradient and discretization error. A critical component of the resampling process entails the reintroduction of random noise to transition the sample $\bm{x}_{t-1}$ back to $\bm{x}_{t}$, effectively resetting the sampling state. The forthcoming lemma substantiates the premise that the act of adding identical noise to two vectors serves to diminish the disparity between the samples.
Lemma 1
[…]
Following Lemma 1, the next proposition demonstrates that the total variation distance between the sampled distribution $q_{t}(\bm{x}_{t}^{i})$, as obtained in the $i$-th resampling of Algorithm 3, and the ground-truth distribution $p_{t}(\bm{x}_{t})$, is progressively reduced through recursive application of the resampling steps.
Proposition 5
Denote $p_{t}(\bm{x}_{t})$ and $q_{t}(\bm{x}_{t})$ as the ground-truth distribution and the distribution distribution sampled by Algorithm 3, and $2\|\bm{x}_{t}\|\leq B$. Then we have
[…]
where $h_{\max}$ is defined in Proposition 3.
The proof is presented in Appendix 8.6. Proposition 5 indicates that the resampling technique serves to “rectify” the distributional divergence caused by guidance. This aspect is crucial, as we have demonstrated that training-free guidance is particularly susceptible to adversarial gradients, which steer the images away from the natural image distribution. The resampling technique realigns these images with the distribution of natural images. This theoretical insight is corroborated by the empirical evidence presented in Figure 2 of [44].
5 Experiments
In this section, we evaluate the efficacy of our proposed techniques across various diffusion models, including CeleA-HQ, ImageNet, and text-to-motion, under different conditions such as sketch, segmentation, text, and object avoidance. We compare our methods with established baselines: Universal Guidance (UG) [2], Loss-Guided Diffusion with Monte Carlo (LGD-MC) [33], and Training-Free Energy-Guided Diffusion Models (FreeDoM) [44]. UG employs guidance as delineated in (5) and uses resampling strategies outlined in Algorithm 3 to enhance performance. Similarly, FreeDoM is founded on (5) and resampling. In addition, FreeDoM incorporates a time-dependent step size for each gradient guidance and judiciously selects the diffusion step for executing resampling. LGD-MC also utilizes guidance from (6), setting the $n=10$. For the sampling process, we employ the DDIM method with 100 steps, consistent with the protocols in FreeDoM and LGD-MC papers. We have developed our approach based on FreeDoM, with the inclusion of random augmentation (Algorithm 1) and projected gradient descent (Algorithm 2).
5.1 Guidance to CelebA-HQ Diffusion
In this subsection, we adopt the experimental setup from [44]. Specifically, we utilize the CelebA-HQ diffusion model [14] to generate high-quality facial images. We explore three guidance conditions: segmentation, sketch, and text. For segmentation guidance, BiSeNet [43] generates the facial segmentation maps, with an $\ell_{2}$-loss applied between the estimated map of the synthesized image and the provided map. Sketch guidance involves using the method from [41] to produce facial sketches, where the loss function is the $\ell_{2}$-loss between the estimated sketch of $\hat{\bm{x}}_{0}$ and the given sketch. For text guidance, we employ CLIP [28] as both the image and text encoders, setting the loss to be the $\ell_{2}$ distance between the image and text embeddings.
For benchmarking, we preprocess the CelebA-HQ dataset with BiSeNet to extract segmentation maps and use the model checkpoint from [41] to acquire the sketches. We randomly select $1000$ samples each of segmentation maps, sketches, and text descriptions. The comparative results are presented in Table 1. Consistent with [44], the resampling number for both FreeDoM and UG is set to $s=1$. Figure 4 displays a random selection of the generated images. More image samples are provided in the supplementary materials.
Methods Segmentation maps Sketches Texts
Distance↓ FID↓ Distance↓ FID↓ Distance↓ FID↓
UG [2] 2247.2 39.91 52.15 47.20 12.08 44.27
LGD-MC [33] 2088.5 38.99 49.46 54.47 11.84 41.74
FreeDoM [44] 1657.0 38.65 34.21 52.18 11.17 46.13
Ours 1575.7 33.31 30.41 41.26 10.72 41.25
Table 1: The performance comparison of various methods on CelebA-HQ with different types of zero-shot guidance. The experimental settings adhere to Table 1 of [44].
5.2 Guidance to ImageNet Diffusion
Methods LGD-MC [33] UG [2] FreeDoM [44] Ours
CLIP Score↑ 24.3 25.7 25.9 27.7
Table 2: The performance comparison of various methods on unconditional ImageNet with zero-shot text guidance. We compare various methods using ImageNet pre-trained diffusion models with CLIP-B/16 guidance. For evaluating performance, the CLIP score is computed using CLIP-L/14.
For the unconditional ImageNet diffusion, we employ text guidance in line with the approach used in FreeDoM and UG [2, 44]. We utilize CLIP-B/16 as the image and text encoder, with cosine similarity serving as the loss function to measure the congruence between the image and text embeddings. To evaluate performance and mitigate the potential for high-scoring adversarial images, we use CLIP-L/14 for computing the CLIP score. In FreeDoM, resampling is conducted for time steps ranging from $800$ to $300$, with the number of resamplings fixed at $10$, as described in [44]. Given that UG resamples at every step, we adjust its resampling number $s=5$ to align the execution time with that of FreeDoM. The textual prompts for our experiments are sourced from [10, 21]. The comparison of different methods is depicted in Table 2. The corresponding randomly selected images are illustrated in Figure 5. The table indicates that our method achieves the highest consistency with the provided prompts. As shown in Figure 5, LGD-MC tends to overlook elements of the prompts. Both UG and FreeDoM occasionally produce poorly shaped animals, likely influenced by adversarial gradients. Our approach addresses this issue through the implementation of random augmentation. Additionally, none of the methods successfully generate images that accurately adhere to positional prompts such as “left to” or “below”. This limitation is inherent to CLIP and extends to all text-to-image generative models [37]. More image samples are provided in the supplementary materials.
5.3 Guidance to Human Motion Diffusion
Methods “Backwards” “Balanced Beam” “Walking” “Jogging”
Loss↓ CLIP↑ Loss↓ CLIP↑ Loss↓ CLIP↑ Loss↓ CLIP↑
Unconditional [36] $3.55+9.66$ 65.6 $47.92+0$ 70.8 $48.88+0$ 37.6 $144.84+0$ 61.72
FreeDoM [44] $1.09+6.63$ 67.23 $9.83+4.48$ 62.65 $1.64+7.55$ 40.12 $34.95+7.83$ 58.74
LGD-MC [33] $0.98+6.48$ 67.31 $4.42+0.02$ 63.13 $1.30+0.39$ 38.82 $6.12+2.38$ 57.89
Ours 0.68+1.32 67.50 1.13+0.30 63.02 0.43+0.31 40.40 2.93+1.15 60.03
Table 3: Comparison of various methods on MDM with zero-shot targeting and object avoidance guidance. Loss is reported as a two-component metric: the first part is the MSE between the target and the actual final position of the individual; the second part measures the object avoidance loss.
In this subsection, we extend our evaluation to human motion generation using the Motion Diffusion Model (MDM) [36], which represents motion through a sequence of joint coordinates and is trained on a large corpus of text-motion pairs with classifier-free guidance. We apply the object avoidance guidance as described in [33]. Let $\bm{x}_{0}(t)$ denote the joint coordinates at time $t$, $\bm{y}_{t}$ the target location, $\bm{y}_{\text{obs}}$ the obstacle location, $r$ the radius of the objects, and $T$ the total number of frames. The loss function is defined as follows:
$\ell=\|\bm{y}_{t}-\bm{x}_{0}(T)\|_{2}^{2}+\sum_{i}\text{sigmoid}(-(\|\bm{x}_{0}(i)-\bm{y}_{\text{obs}}\|-r)\times 50)\times 100.$ (9)
Our experimental configuration adheres to the guidelines set forth in [33]. We assess the methods using the targeting loss (the first term in (9)), the object avoidance loss (the second term in (9)), and the CLIP score calculated by MotionCLIP [35]. For FreeDoM, the number of resamplings is set to $3$. In our method, random augmentation is omitted because the guidance is not computed by neural networks so the adversarial gradient issues are absent. The quantitative results of our investigation are summarized in Table 3, while Figure 6 showcases randomly selected samples. Our methods exhibit enhanced control quality over the generated motion. The videos are provided in the supplementary materials.
6 Conclusions
In this paper, we conducted a comprehensive investigation into training-free guidance, which employs pre-trained diffusion models and guides them using the off-the-shelf trained on clean images. Our exploration delved into the underlying mechanisms and fundamental limits of these models. Moreover, we proposed a set of enhancement techniques and verified their effectiveness both theoretically and empirically.
Limitations. Despite our efforts to mitigate the shortcomings of training-free methods and enhance their performance, certain intractable limitations remain. Notably, the refined training-free guidance still necessitates a higher number of NFEs when compared with extensive training methods such as classifier-free guidance. This is because adversarial gradient cannot be fully eliminated without training.
Ethical Consideration. Similar to other models designed for image creation, our model also has the unfortunate potential to be used for creating deceitful or damaging material. We pledge to restrict the usage of our model exclusively to the realm of research to prevent such misuse.
Figure 4: Qualitative results of CelebA-HQ with zero-shot segmentation, sketch, and text guidance. The images are randomly selected.
Figure 5: Qualitative results of ImageNet model with zero-shot text guidance. The images are randomly selected.
Figure 6: Qualitative results of human motion diffusion with zero-shot object avoidance and targeting guidance. Instances of intersection with obstacles are highlighted by marking the person in red. The trajectories are randomly selected.